# When Radiation Meets Linux: Analyzing Soft Errors in Linux on COTS SoCs under Proton Irradiation

### v2

@@ Abstract
In space, radiation exposure in COTS electronics degrades electronic reliability through three primary phenomena: single-event effects (SEEs), total ionizing dose (TID), and displacement damage (DD). While TID and DD accumulate over time through distinct physical mechanisms, SEEs occur when energetic particles like protons, heavy ions, or neutrons via direct or indirect ionization, deposit charges after penetrating a semiconductor substrate, potentially inducing transient soft errors or permanent hard errors. Soft errors occur when the collected charge () at a circuit’s sensitive node exceeds the critical charge threshold (), inducing transient faults that either change the output state or cause a transient spike that gradually dissipates. Unlike hard errors, which may cause permanent silicon damage, including unmitigated latch-up, soft errors are transient, nondestructive, and repairable. Soft errors, central to our study, include single-event upsets (SEUs) and single-event functional interrupts (SEFIs), which are:
@@ II-B Stress Testing Tool: Workload and Test Parameters
We employed the stress-ng utility [45], a widely used software stress tool, to expose vulnerable hardware and software components likely to fail under combined workload stress and radiation-induced failures. Using the stress-ng tool, we exercised all four CPU cores of the ARM Cortex-A53 SoC, including critical SoC components, such as I/O peripherals, virtual memory (VM) and hard disk drive (HDD) storage. We used Fast Fourier Transforms (FFTs) as the choice of workload for computational stress on these SoCs, primarily due to the highly memory-bound nature of FFTs. Operationally, FFTs decompose a recursive or iterative data set into smaller components, requiring frequent memory access to move data between memory and processing units. We anticipated that intensive FFT workloads operating on corrupted data affected by soft errors could lead to elevated fault propagation. In Table V, Table VI, and Table VII, we present the Linux error data from the irradiation test, showcasing the effectiveness of our unique, exhaustive test parameters in capturing a diverse range of soft errors not reported in similar studies.
@@ II-E DUT Selection
Our DUT selection, particularly SoCs, prioritized SWaP attributes and cost-effectiveness for CubeSat and NanoSat missions, targeting LEOs with relatively low radiation levels compared to higher orbits. Additional factors, such as long-term product support availability (e.g., >5 years lifecycle) and an active development community, were also prioritized during DUT selection. Our test objectives required SoCs capable of running 32 or 64-bit Linux with adequate system resources. After evaluating all criteria, we selected the following devices: i) Raspberry Pi Zero 2 W, SoC on a Single-Board Computer (SBC); ii) NXP i.MX 8M Plus, SoC on a System-on-Module (SoM); and iii) OrangeCrab v2 ECP5 FPGA, SoC on a RISC-V softcore—all capable of running the Linux OS. In Table I, we briefly compare the hardware and software specifications of all three DUTs.
@@ II-E1 DUT 1: Raspberry Pi Zero 2 W
Software: The rpi02w operated on the vendor-supplied Raspbian Linux variant, configured with minimal adjustments to streamline testing and logging: (i) The Pi was set to auto-login as root on every reboot, ensuring testing tools ran with root privileges to prevent permission issues during tests. (ii) We ensured test logs were directed to an auto-mounting external USB flash drive, including real-time logging from the external monitoring unit. However, if the auto-mount of the USB fails at boot time, it can lead to a longer "boot delay," until a proper timeout is defined. A notable limitation of Pi’s is the absence of an onboard Real-Time Clock (RTC), which tracks the current time and date, even when the system is powered off. This causes the system to lose track of time at startup if no time synchronization is configured from the NTP servers. If not configured, it complicates the timestamping of the log files critical for post-processing.
@@
Software: The i.MX 8M plus board has native support for Portainer [52], a container management tool designed to deploy and manage containers, including Docker images. We packaged our complete testing software suite for the irradiation test within the Debian-based Docker image and used it during the irradiation test.
@@ III-C1 DUT 1:Raspberry Pi Zero 2 W (40 nm CMOS)
DUT 2 consistently exhibited the lowest Linux-SEFI cross-section, 5–14x lower than DUT 1 and 3, as detailed in Table III(b). Compared to DUT 1, which also features the ARM Cortex-A53 SoC, significant SEE cross-section variation was observed. For instance, the highest was observed at 30 MeV () and lowest at 20 MeV (), i.e. lower than DUT 1’s lowest (Observed) cm2 at 58 MeV. Notably, 3 out of 37 Linux test runs completed without crashes in their intended beam run-time, as shown in Table II(b). The i.MX 8M Plus SoC, built on a 14 nm Low-Power Compact (LPC) FinFET process, is known for improved current control and reduced charge leakage with its 3D gate structure to bulk CMOS [62]. During irradiation, the proton beam targeted the SoC (from the top) and 4 GB LPDDR4 module. The SoC integrates a GPU, NPU, connectivity interfaces, and various I/O peripherals. Unlike SiPs, SoMs arrange components horizontally onto a shared PCB substrate, ensuring a relatively uniform exposure with the incident radiation beam.
@@ III-C3 DUT 3: OrangeCrab v2 FPGA (40 nm CMOS)
In Table IV, we provide the estimated Linux crash rate for DUT2, targeting the ISS orbit at 500 km and 51.6 degrees using the CREME96 tool [63]. The analysis includes a pre-calculated trapped proton model (TRP) for both solar minimum and maximum cycles, the Geomagnetic Transmission Function (GTRN) under Quiet and Stormy conditions, and 100 mils of aluminum shielding with the Nuclear Transport Routine (TRANS) module. A Weibull fit was estimated for the SEE cross-section, as the Weibull function more accurately models proton-induced SEE cross-sections compared to the Bendel parameters used in the CREME96 tool, as noted in [64]. Our estimated Linux crash rates align with Millet et al. [3], whose analysis also reveals the paradoxical decrease rate of reboots per year during Solar Max cycles. This phenomenon is attributed to the Sun’s stronger magnetic field and solar wind during solar max, which significantly enhances heliospheric shielding. As a result, Galactic Cosmic Rays (GCRs) are effectively deflected, reducing the particle flux responsible for soft errors.
| Linux Crashes Per Day | Linux Crashes Per Year | |||
| SoC | Solar Min. (Avg-Quiet) | Solar Max. (Avg-Stormy) | Solar Min. (Avg-Quiet) | Solar Max. (Avg-Stormy) |
| imx8mp | 2.13E-03 | 1.20E-03 | 0.78 | 0.44 |
@@ IV-1 DUT 1: Raspberry Pi Zero 2 W
Our experimental data on Linux errors, outlined in Table V, reports frequent low-level MMU corruptions, cascading failures across multimedia (vc4-drm), Bluetooth, firmware transaction timeouts, and memory errors to DUT 2 and 3. Under irradiation, DUT 1 experienced multiple reboot failures, likely due to disruptions in the boot process, including boot file corruption and power-related issues. Additionally, overcurrent protection (OCP) events were logged by the external PPS (power supply) unit, indicating that DUT 1’s power rails are more sensitive to high-current incidents, possibly latch-ups, compared to DUT 2 and DUT 3. The DUT recovered after a soft reboot; however, most of the time, it failed to reboot under the beam. A total of 43% of Linux errors were classified as critical (non-recoverable), affecting core computational and power control functions. These included kernel oops, kernel paging errors, and firmware request failures.
@@
Low-impact errors represented 28% of all cases. They indicate errors with minimal reliability impact, including kernel taints and stress-ng exposing row hammer, modulo-X, and gray code errors under irradiation. These faults had minimal effect on overall DUTs’ reliability during irradiation. Overall, 78% of all detected Linux errors were attributed to the memory management (mm) and device driver (dd) subsystems, as illustrated in pie chart 8. These findings indicate that the CPU and memory subsystems were the most vulnerable components in DUT 1. These susceptible areas in the Linux kernel should be prioritized for software-based mitigation strategies, as proposed in Table V.

### v3

@@ Abstract
In space, radiation exposure in COTS electronics degrades electronic reliability through three primary phenomena: single-event effects (SEEs), total ionizing dose (TID), and displacement damage (DD). While TID and DD accumulate over time through distinct physical mechanisms, SEEs occur when energetic particles like protons, heavy ions, or neutrons, via direct or indirect ionization, deposit charges after penetrating a semiconductor substrate, potentially inducing transient soft errors or permanent hard errors. Soft errors occur when the collected charge () at a circuit’s sensitive node exceeds the critical charge threshold (), inducing transient faults that either change the output state or cause a transient spike that gradually dissipates. Unlike hard errors, which may cause permanent silicon damage, including unmitigated latch-up, soft errors are transient, nondestructive, and repairable. Soft errors, central to our study, include single-event upsets (SEUs) and single-event functional interrupts (SEFIs), which are:
@@ II-B Stress Testing Tool: Workload and Test Parameters
We employed the stress-ng utility [45], a widely used software stress tool, to expose vulnerable hardware and software components likely to fail under combined workload stress and radiation-induced failures. Using the stress-ng tool without Linux root privileges, we exercised all four CPU cores of the ARM Cortex-A53 SoC, including critical SoC components such as I/O peripherals, virtual memory (VM), and hard disk drive (HDD) storage. We used Fast Fourier Transforms (FFTs) as the choice of workload for computational stress on these SoCs, primarily due to the highly memory-bound nature of FFTs. Operationally, FFTs decompose a recursive or iterative data set into smaller components, requiring frequent memory access to move data between memory and processing units. We anticipated that intensive FFT workloads operating on corrupted data affected by soft errors could lead to elevated fault propagation. In Table V, Table VI, and Table VII, we present the Linux error data from the irradiation test, showcasing the effectiveness of our unique, exhaustive test parameters in capturing a diverse range of soft errors not reported in similar studies.
@@ II-E DUT Selection
Our DUT selection, particularly SoCs, prioritized SWaP attributes and cost-effectiveness for CubeSat and NanoSat missions, targeting LEOs with relatively low radiation levels compared to higher orbits. Additional factors, such as long-term product support availability (e.g., >5 years lifecycle) and an active development community, were also prioritized during DUT selection. Our test objectives required SoCs capable of running 32 or 64-bit Linux with adequate system resources. After evaluating all criteria, we selected the following devices: i) Raspberry Pi Zero 2 W, SoC on a Single-Board Computer (SBC); ii) NXP i.MX 8M Plus, SoC on a System-on-Module (SoM); and iii) OrangeCrab v2 ECP5 FPGA, SoC on a RISC-V softcore, all capable of running the Linux OS. In Table I, we briefly compare the hardware and software specifications of all three DUTs.
@@ II-E1 DUT 1: Raspberry Pi Zero 2 W
Software: The rpi02w operated on the vendor-supplied Raspbian Linux variant, configured with minimal adjustments to streamline testing and logging: (i) The Pi was set to auto-login as root on every reboot, ensuring testing tools ran with root privileges to prevent permission issues during tests. (ii) We ensured test logs were directed to an auto-mounting external USB flash drive, including real-time logging from the external monitoring unit. However, if the auto-mount of the USB fails at boot time, it can lead to a longer "boot delay" until a proper timeout is defined. A notable limitation of Pi’s is the absence of an onboard Real-Time Clock (RTC), which tracks the current time and date, even when the system is powered off. This causes the system to lose track of time at startup if no time synchronization is configured from the NTP servers. If not configured, it complicates the timestamping of the log files critical for post-processing.
@@
Software: The i.MX 8M Plus board natively supports Portainer [52] for managing Docker images. In our test, stress-ng ran in a non-privileged Docker container, which, unlike virtual machines (VMs), uses the host Linux kernel for its functioning. Docker also uses settings like namespaces and cgroups to manage processes’ requirements for CPU, memory, and I/O from the underlying hardware. However, despite Docker’s process isolation, data show containerized applications perform identically to native (Linux) process execution, aside from a few seconds’ initialization delays from the Docker daemon [66]. Likewise, Docker processes share the same memory management and Linux kernel’s default Address Space Layout Randomization (ASLR) memory assignment for its processes as host user-space Linux processes. Thus, despite the isolation, the memory used by processes running within Docker remains equally exposed to SEEs unless it is explicitly confined to a specific memory region.
@@ III-C1 DUT 1:Raspberry Pi Zero 2 W (40 nm CMOS)
DUT 2 consistently exhibited the lowest Linux-SEFI cross-section, 5–14x lower than DUT 1 (which also features the ARM Cortex-A53 SoC) and DUT 3, as detailed in Table III(b). For instance, the highest was observed at 30 MeV () and lowest at 20 MeV (), i.e. lower than DUT 1’s lowest (Observed) cm2 at 58 MeV. Notably, 3 out of 37 Linux test runs completed without crashes in their intended beam run-time, as shown in Table II(b). The i.MX 8M Plus SoC, built on a 14 nm Low-Power Compact (LPC) FinFET process, is known for improved current control and reduced charge leakage with its 3D gate structure to bulk CMOS [63]. During irradiation, the proton beam targeted the SoC (from the top) and 4 GB LPDDR4 module. The SoC integrates a GPU, NPU, connectivity interfaces, and various I/O peripherals. Unlike SiPs, SoMs arrange components horizontally onto a shared PCB substrate, ensuring a relatively uniform exposure to the incident radiation beam. A Weibull fit for this DUT was estimated from its – range, as shown in Table IV.
@@ III-C3 DUT 3: OrangeCrab v2 FPGA (40 nm CMOS)
| (MeV) | (MeV) | |||
| 20 | 3.37 | 2 |
@@ IV-1 DUT 1: Raspberry Pi Zero 2 W
Our experimental data on Linux errors, outlined in Table V, reports frequent low-level MMU corruptions, cascading failures across multimedia (vc4-drm), Bluetooth, firmware transaction timeouts, and memory errors to DUT 2 and 3. Under irradiation, DUT 1 experienced multiple reboot failures, likely due to disruptions in the boot process, including boot file corruption and power-related issues. Additionally, overcurrent protection (OCP) events were logged by the external PPS (power supply) unit, indicating that DUT 1’s power rails are more sensitive to high-current incidents, possibly micro latch-ups. The DUT recovered after a soft reboot; however, most of the time, it failed to reboot under the beam. A total of 43% of Linux errors were classified as critical (non-recoverable), affecting core computational and power control functions. These included kernel oops, kernel paging errors, and firmware request failures.
@@
Low-impact errors represented 28% of all cases. They indicate errors with minimal reliability impact, including kernel taints. These faults had minimal effect on overall DUTs’ reliability during irradiation. Overall, 78% of all detected Linux errors were attributed to the memory management (mm) and device driver (dd) subsystems, as illustrated in pie chart 8. The susceptibility of Linux drivers to frequent crashes aligns with prior studies showing their error rate is three to seven times higher than the rest of the kernel [54]. Jointly, these findings indicate that the CPU and memory subsystems were the most vulnerable components in DUT 1. These susceptible areas in the Linux kernel should be prioritized for software-based mitigation strategies, as proposed in Table V.